OpenStudy (anonymous):

Find the equation of the line with slope 2/3 and y-intercept 4

OpenStudy (anonymous):

y=mx+C where m=2/3 and c=4

OpenStudy (anonymous):

y = 2/3 x + 4

OpenStudy (anonymous):

it should say "write" rather than "find" because you are given the information you need to write it down